A hot-air balloon rises at a velocity of 10 m/s. When the balloon is 60 meters above the ground, it releases a sandbag. After release, the sandbag is in free fall. Using the four kinematics equations listed below, provide your simplified equations using a coordinate system whose y axis is positive up and the x axis positive right such that the origin is located at the ground level

tion: x=xo + vt 2 22a(-
Answer #1

Here, the motion is entirely vertical and so x = xo = 0 m

u = uy = initial velocity of the sandbag = 10 m/s

yo = 60 m and a = g = - 9.8 m/s2

so, y = 60 + 10t - (1/2)gt2 = 60 + 10t - 4.9t2

v = 10 - 9.8t

and v2 = 100 - 19.6(y - 60)

these are the equations of motion for the sandbag.